# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@ 2.1 Differential equations as submanifolds of jet spaces
[…]
A $k$-th order system of differential equations $\mathcal{E}$, for the sections of $\pi$, can be geometrically interpreted as a submanifold $\mathcal{E}\subset J^{k}(\pi)$ and its solutions are just sections $s$ of $\pi$ whose $k$-order prolongations $j_{k}(s)$ lay on $\mathcal{E}$. Hence, solutions of $\mathcal{E}$ are sections of $\pi$ whose $k$-order prolongation are integral manifolds of the induced Cartan distribution $\mathcal{C}^{k}(\mathcal{E}):=\mathcal{C}^{k}(\pi)\cap T\mathcal{E}$ over $\mathcal{E}$.
[…]
Also $J^{\infty}(\pi)$ is naturally equipped with a Cartan distribution denoted by $\mathcal{C}(\pi)$ and defined as the inverse limit of the sequence of surjections $\mathcal{C}^{1}(\pi)\overset{\pi_{1,0*}}{\longleftarrow}\mathcal{C}^{2}(\pi)\overset{\pi_{2,1*}}{\longleftarrow}...\overset{\pi_{k,k-1*}}{\longleftarrow}\mathcal{C}^{k}(\pi)\overset{\pi_{k+1,k*}}{\longleftarrow}...\;.$ In coordinates $\mathcal{C}(\pi)$ can be described either as the annihilator of the Pfaffian system $\{\omega_{\sigma}^{j}=du_{\sigma}^{j}-\sum_{i}u_{\sigma+1_{i}}^{j}dx_{i}:\;|\sigma|\geq 0,\;j=1,...,m\}$, or as the distribution generated by the totality of vector fields $\{\partial_{u_{\sigma+1_{i}}^{j}},\,D_{i}:\;|\sigma|\geq 0,\;j=1,...,m\}$, where
$D_{i}:=\partial_{x_{i}}+{\displaystyle\sum_{|\rho|\geq 0}\sum_{j=1}^{m}u_{\rho+1_{i}}^{j}\partial_{u_{\rho}^{j}}}$
denote the total derivative operators.
[…]
By definition the infinite prolongation ()=$\mathcal{E}^{(\infty)}=\{D_{\mu}\mathbf{F}=\mathbf{0}:\;|\mu|\geq 0\}$, of a formally integrable equation $\mathcal{E}$, is the inverse limit of the sequence of fiber bundles $\pi_{k+r+1,k+r}:\mathcal{E}^{(r+1)}\rightarrow\mathcal{E}^{(r)}$. By restricting $\Lambda^{*}(\pi)$ to $\mathcal{E}^{(\infty)}$ one gets the exterior algebra $\Lambda^{*}(\mathcal{E})$ of differential forms on $\mathcal{E}^{(\infty)}$ and in particular the algebra $\mathcal{F}(\pi)$ of smooth functions on $\mathcal{E}^{(\infty)}$. Moreover, since any prolongation $\mathcal{E}^{(r)}$ is naturally equipped with an induced Cartan distribution $\mathcal{C}^{k+r}(\mathcal{E}^{(r)})$, also $\mathcal{E}^{(\infty)}$ is naturally equipped with an induced Cartan distribution denoted by $\mathcal{C}(\mathcal{E})$ and defined as the inverse limit of the sequence of surjections $\mathcal{C}^{k}(\mathcal{E})\overset{\pi_{k+1,k*}}{\longleftarrow}\mathcal{C}^{k+1}(\mathcal{E}^{(1)})\overset{\pi_{k+2,k+1*}}{\longleftarrow}...\;$. In coordinates, one has that $\mathcal{C}(\mathcal{E})=\left\langle\bar{D}_{1},...,\bar{D}_{n}\right\rangle=\text{Ann}\left\{\bar{\omega}_{\sigma}^{j}:\,j=1,...,m,\quad|\sigma|\geq 0\right\}$, where $\bar{D}_{i}$ and $\bar{\omega}_{\sigma}^{j}$ are the restrictions to $\mathcal{E}^{(\infty)}$ of the total derivatives and Cartan forms, respectively.
Now, since $\mathcal{C}(\pi)$ is totally horizontal with respect to the mapping $\pi_{\infty}:J^{\infty}(\pi)\rightarrow M$, the tangent bundle $\mathcal{T}\left(\pi\right)$ on $J^{\infty}(\pi)$ decomposes as $\mathcal{T}\left(\pi\right)=\mathcal{V}(\pi)\oplus\mathcal{C}(\pi)$, where $\mathcal{V}(\pi):=Ker\left(\pi_{\infty}\right)_{*}$. Dually one has $\Lambda^{1}(\pi)=\Lambda^{(1,0)}(\pi)\oplus\Lambda^{(0,1)}(\pi)$, where $\Lambda^{(1,0)}(\pi):=Ann\left(\mathcal{V}(\pi)\right)$ and $\Lambda^{(0,1)}(\mathcal{\pi}):=Ann\left(\mathcal{C}(\pi)\right)$ are the $\mathcal{F}(\pi)$-modules of horizontal and vertical $1$-forms on $J^{\infty}(\pi)$ locally generated by $\{dx_{i}\}$ and Cartan forms $\left\{\omega_{\sigma}^{j}\right\}$, respectively. More in general, by considering $\Lambda^{(p,q)}(\pi)=\left(\bigwedge^{p}\Lambda^{(1,0)}(\pi)\right)\bigwedge\left(\bigwedge^{q}\Lambda^{(0,1)}(\pi)\right)$, the $\mathcal{F}(\pi)$-module of $r$-forms on $J^{\infty}(\pi)$ decomposes as $\Lambda^{r}\left(\pi\right)=\bigoplus_{p+q=r}\Lambda^{(p,q)}\left(\pi\right)$. By definition we set $\mathcal{F}(\pi)=\Lambda^{(0,0)}(\pi)$. Accordingly, the exterior differential splits into the sum $d=d_{H}+d_{V}$ of the horizontal and vertical differentials $d_{H}:\Lambda^{(p,q)}(\mathcal{\pi})\rightarrow\Lambda^{(p+1,q)}(\pi)$ and $d_{V}:\Lambda^{(p,q)}(\pi)\rightarrow\Lambda^{(p,q+1)}(\pi)$, satisfying $d_{H}^{2}=d_{V}^{2}=0$ and $d_{H}\circ d_{V}=-d_{V}\circ d_{H}$. In coordinates, horizontal and vertical differentials can be easily computed, since on any function $f\in\mathcal{F}(\pi)$ they act as graded derivations on $\Lambda^{*}(\pi)$ such that $d_{H}f:=\sum_{i}D_{i}(f)dx_{i}$ and $d_{V}f:=\sum_{\sigma}\sum_{j}\frac{\partial f}{\partial u_{\sigma}^{j}}\omega_{\sigma}^{j}$.
Analogously, given a formally integrable equation $\mathcal{E}$, since $\mathcal{C}(\mathcal{E})$ is totally horizontal with respect to the mapping $\bar{\pi}_{\infty}:\mathcal{E}^{(\infty)}\rightarrow M$, the tangent bundle $\mathcal{T}\left(\mathcal{E}\right)$ on $\mathcal{E}^{(\infty)}$ decomposes as $\mathcal{T}\left(\mathcal{E}\right)=\mathcal{V}(\mathcal{E})\oplus\mathcal{C}(\mathcal{E})$, where $\mathcal{V}(\mathcal{E}):=Ker\left(\bar{\pi}_{\infty*}\right)$ is the vertical bundle on $\mathcal{E}^{(\infty)}$. Hence the $\mathcal{F}(\mathcal{E})$-modules $\Lambda^{(1,0)}(\mathcal{E})$ and $\Lambda^{(0,1)}(\mathcal{E})$ of horizontal and vertical $1$-forms on $\mathcal{E}^{(\infty)}$, locally generated by $\{dx^{i}\}$ and restricted Cartan forms $\left\{\bar{\omega}_{\sigma}^{j}:=\left.\omega_{\sigma}^{j}\right|_{\mathcal{E}^{(\infty)}}\right\}$, can be used to decompose the $\mathcal{F}(\mathcal{E})$-module of $r$-forms on $\mathcal{E}^{(\infty)}$ as $\Lambda^{r}\left(\mathcal{E}^{(\infty)}\right)=\bigoplus_{p+q=r}\Lambda^{(p,q)}\left(\mathcal{E}\right)$. By definition we set $\mathcal{F}(\mathcal{E})=\Lambda^{(0,0)}(\mathcal{E})$. Accordingly, on $\mathcal{E}^{(\infty)}$ the exterior differential $d$ (still denoted by $d$, for ease of notation) splits into the sum $d=\bar{d}_{H}+\bar{d}_{V}$ of the horizontal and vertical differentials $\bar{d}_{H}:\Lambda^{(p,q)}(\mathcal{E})\rightarrow\Lambda^{(p+1,q)}(\mathcal{E})$ and $\bar{d}_{V}:\Lambda^{(p,q)}(\mathcal{E})\rightarrow\Lambda^{(p,q+1)}(\mathcal{E})$, which satisfy $\bar{d}_{H}^{2}=\bar{d}_{V}^{2}=0$ and $\bar{d}_{H}\circ\bar{d}_{V}=-\bar{d}_{V}\circ\bar{d}_{H}$. In coordinates, $\bar{d}_{H}$ and $\bar{d}_{V}$ can be easily computed, since they act as graded derivations on $\Lambda^{*}(\mathcal{E}^{(\infty)})$ and for any function $f\in\mathcal{F}(\mathcal{E})$ they are such that $\bar{d}_{H}f:=\sum_{i}\bar{D}_{i}(f)dx_{i}$ and $\bar{d}_{V}f:=\sum_{\sigma}\sum_{j}\frac{\partial f}{\partial u_{\sigma}^{j}}\bar{\omega}_{\sigma}^{j}$, where $\bar{D}_{i}$ denote the total derivatives restricted to $\mathcal{E}^{(\infty)}$. For ease of notation, we will denote $\Lambda^{(p,0)}(\pi)$ and $\Lambda^{(p,0)}(\mathcal{E})$ by $\bar{\Lambda}^{p}(\pi)$ and $\bar{\Lambda}^{p}(\mathcal{E})$, respectively.
Starting from these algebraic structures in the algebra of differential forms on $\mathcal{E}$, new constructions and new notions can be introduced further. For instance, using the horizontal differentials $\bar{d}_{H}:\overline{\Lambda}^{p}(\mathcal{E})\rightarrow\overline{\Lambda}^{p+1}(\mathcal{E})$, one can consider the notion of conservation law for $\mathcal{E}$, which is a closed horizontal form $\mu\in\overline{\Lambda}^{n-1}(\mathcal{E})$, i.e., a horizontal $(n-1)$-form on $\mathcal{E}^{(\infty)}$ satisfying $\bar{d}_{H}\mu=0$. On the other hand, since an exact horizontal form is trivially closed, one can naturally limit itself to consider closed horizontal $(n-1)$-forms up to exact horizontal forms. Thus, a conservation law can also be understood as a cohomology class $[\mu]\in\overline{H}^{n-1}(\mathcal{E})$, i.e., $[\mu]=\left\{\mu+\bar{d}_{H}\rho:\,\rho\in\overline{\Lambda}^{n-2}(\mathcal{E})\right\}$. For instance, when $n=2$, a conservation law is locally described by an horizontal $1$-form $Adx_{1}+Bdx_{2}$ on $\mathcal{E}^{(\infty)}$ such that $\bar{D}_{2}A-\bar{D}_{1}B=0$, i.e., $D_{2}A-D_{1}B=0$ on $\mathcal{E}^{(\infty)}$.
Moreover, given a matrix Lie algebra $\mathfrak{g}$, one may consider the exterior algebras $\mathfrak{g}\otimes\Lambda^{*}(\pi)$ and $\mathfrak{g}\otimes\Lambda^{*}(\mathcal{E}^{(\infty)})$ of $\mathfrak{g}$-valued forms on $J^{\infty}(\pi)$ and $\mathcal{E}^{(\infty)}$, respectively. The graded algebra of $\mathfrak{g}$-valued horizontal forms on $J^{\infty}(\pi)$ and $\mathcal{E}^{(\infty)}$, will be denoted by $\mathfrak{g}\otimes\bar{\Lambda}^{*}\left(\pi\right)=\bigoplus_{p}\mathfrak{g}\otimes\bar{\Lambda}^{p}(\pi)$ and $\mathfrak{g}\otimes\bar{\Lambda}^{*}\left({\cal E}\right)=\bigoplus_{p}\mathfrak{g}\otimes\bar{\Lambda}^{p}\left({\cal E}\right)$, respectively. By definition, $\mathfrak{g}$-valued horizontal $p$-forms on $J^{\infty}(\pi)$ (resp., $\mathcal{E}^{(\infty)}$) are generated by $\mathfrak{g}$-valued $p$-forms $A\omega$, with $A$ $\mathfrak{g}$-valued functions on $J^{\infty}(\pi)$ (resp., $\mathcal{E}^{(\infty)}$). Then, one may define a skew-symmetric product [,][\;,\;][ , ] by linearly extending the product $\left[A_{1}\omega_{1},A_{2}\omega_{2}\right]:=\left[A_{1},A_{2}\right]\omega_{1}\wedge\omega_{2}$, between generators. One can check that [,][\;,\;][ , ] satisfies the following properties: (i) $[\rho,\sigma]=-(-1)^{rs}[\sigma,\rho]$; (ii) $(-1)^{rt}[\rho,[\sigma,\tau]]+(-1)^{sr}[\sigma,[\tau,\rho]]+(-1)^{ts}[\tau,[\rho,\sigma]]=0$; (iii) $d_{H}[\rho,\sigma]=[d_{H}\rho,\sigma]+(-1)^{r}[\rho,d_{H}\sigma]$, analogously for $\bar{d}_{H}$. Where $r,s$ and $t$ are the degrees of $\rho,\sigma$ and $\tau$, respectively. Also, one may define an exterior product $\wedge$ on $\mathfrak{g}\otimes\bar{\Lambda}^{*}\left(\pi\right)$, or $\mathfrak{g}\otimes\bar{\Lambda}^{*}\left(\mathcal{E}\right)$, by linearly extending the product $A_{1}\omega_{1}\wedge A_{2}\omega_{2}=A_{1}A_{2}\omega_{1}\wedge\omega_{1}$.
This allows one to introduce the notion of zero-curvature representation (ZCR) of ${\cal E}$, that is a $\mathfrak{g}$-valued non-vanishing $1$-form $\alpha\in\mathfrak{g}\otimes\bar{\Lambda}^{1}\left({\cal E}\right)$ such that
@@
This is a very important notion in the theory of integrable equations with $2$ independent variables [1, 41, 44]. In such a case, by taking $\alpha=A\,dx_{1}+Bdx_{2}$, condition (1) reads $D_{2}A-D_{1}B+[A,B]=0$ on $\mathcal{E}^{(\infty)}$. Here we notice that (1) can also be equivalently written as $\bar{d}_{H}\alpha-\alpha\wedge\alpha=0$. Moreover, since $\mathcal{E}^{(\infty)}\subset J^{\infty}(\pi)$, any element of $\mathfrak{g}\otimes\bar{\Lambda}^{1}(\mathcal{E})$ can be identified with an element of $\mathfrak{g}\otimes\bar{\Lambda}^{1}(\pi)$. Hence, in the outer geometry, (1) can also be rewritten as $d_{H}\alpha-\frac{1}{2}\left[\alpha,\alpha\right]=0\;\mbox{mod}\,\mathcal{E}^{(\infty)}$. In particular, when =$\mathcal{E}=\{F^{j}=0,\,j=1,...,h\}$, under regularity assumptions (i.e., if any prolongation $\mathcal{E}^{(\infty)}$, $h\geq 0$, is totally non-degenerating [34]) equation (1) can also be rewritten in “characteristic” form $d_{H}\alpha-\frac{1}{2}\left[\alpha,\alpha\right]=\sum D_{\sigma}(F^{j})\omega_{j}^{\sigma}$, where $\omega_{j}^{\sigma}\in\mathfrak{g}\otimes\bar{\Lambda}^{2}(\pi)$. Hence in general (1) holds only modulo some differential consequences of $\mathcal{E}$, thus in general (1) is not equivalent to $\{F^{j}=0,\,j=1,...,h\}$.
@@ 2.2 $\mathcal{C}$-morphisms (or Lie-Bäcklund maps)
A $\mathcal{C}$-morphism, also referred to as a Lie-Bäcklund transformations [4, 22], is a smooth map $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$, such that $\mathcal{B}_{*}\mathcal{C}_{\theta}(\pi)\subset\mathcal{C}_{\mathcal{B}(\theta)}(\pi^{\prime})$, for any $\theta\in J^{\infty}(\pi)$, i.e.,
$\mathcal{B}_{*}\mathcal{C}(\pi)\subset\mathcal{C}(\pi^{\prime}).$ (2)
Notice that $\mathcal{B}_{*}\mathcal{C}(\pi)\subset\mathcal{C}(\pi^{\prime})$ is equivalent to $\mathcal{B}^{*}\left(\mathcal{I}_{\mathcal{C}(\pi^{\prime})}\right)\subset\mathcal{I}_{\mathcal{C}(\pi)}$, where $\mathcal{I}_{\mathcal{C}(\pi)}$ is the EDS generated by Cartan (or multi-contact) forms $\omega_{\sigma}^{j}=du_{\sigma}^{j}-\sum_{i}u_{\sigma+1_{i}}^{j}dx_{i}$ on $J^{\infty}(\pi)$, analogously $\mathcal{I}_{\mathcal{C}(\pi^{\prime})}$.
Notice that in general these transformations are generalizations of Lie transformations (i.e., point and contact transformations) that need not to be diffeomorphisms. Indeed, as shown by Bäcklund [6] (see also [4]), only Lie transformations are invertible. Moreover, by $\mathcal{B}^{*}\left(\mathcal{I}_{\mathcal{C}(\pi^{\prime})}\right)\subset\mathcal{I}_{\mathcal{C}(\pi)}$ it follows that under such a transformation the image of an integral manifolds of Cartan distribution is still an integral manifold.
@@
$\left\{\begin{array}[]{l}x^{\prime}_{i}=\xi^{i}\left(x,u^{(k)}\right),\vspace{10pt}\\ u^{\prime}{}^{j}=\nu^{j}\left(x,u^{(k)}\right),\vspace{10pt}\\ \vdots\\ u^{\prime}{}_{\sigma}^{j}=\nu_{\sigma}^{j}\left(x,u^{(k+|\sigma|)}\right),\qquad|\sigma|\geq 0\end{array}\right.$ (3)
where in view of $\mathcal{B}^{*}\left(\mathcal{I}_{\mathcal{C}(\pi^{\prime})}\right)\subset\mathcal{I}_{\mathcal{C}(\pi)}$ the functions $\xi^{i}$ and $\nu_{\sigma}^{j}$ are such that $d\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d\xi^{i}=0\mod{\mathcal{I}_{\mathcal{C}}}$.
[…]
@@
?proofname? .
Indeed, being $\Sigma$ the infinite prolongation of a section $\mathfrak{s}$ locally described as $\mathfrak{s}(x)=(x,u(x))$, one has $\left(\mathcal{B}\circ\mathfrak{s}^{(\infty)}\right)^{*}(\omega^{\prime}{}_{\sigma}^{j})=\allowbreak\left(\mathfrak{s}^{(\infty)}\right)^{*}\circ\mathcal{B}^{*}\left(\omega^{\prime}{}_{\sigma}^{j}\right)\allowbreak=0\mod{\mathcal{I}_{\mathcal{C}(\pi)}}$, since the graph of $\mathfrak{s}^{(\infty)}$ is an integral manifold of $\mathcal{I}_{\mathcal{C}(\pi)}$. Thus the graph $\Sigma^{\prime}$ of $\mathcal{B}\circ\mathfrak{s}^{(\infty)}$describes another integral manifold of $\mathcal{I}_{\mathcal{C}(\pi^{\prime})}$. On the other hand whenever $\det\left(D_{s}\xi^{i}\right)\neq 0$ along the graph of $\mathfrak{s}^{(k)}$ (i.e., $\mathfrak{s}^{(k)}{}^{*}\left(\det\left(D_{s}\xi^{i}\right)\right)\neq 0$), the first $n$-equations $\{x^{\prime}_{i}=\xi^{i}(x,u^{(k)}(x)),\,i=1,...,n\}$ of (3) restricted to $\mathfrak{s}^{(k)}$ can be locally solved with respect to $(x_{1},...,x_{n})$. Thus, there exists a local diffeomorphism $x^{\prime}=\mathcal{B}_{\mathfrak{s}}(x)=\left(\xi\circ\mathfrak{s}^{(k)}\right)(x)$ that allows one to pass from the parametrization $\left(\mathcal{B}\circ\mathfrak{s}^{(\infty)}\right)(x)$ of $\Sigma^{\prime}$ to the new parametrization$\left(\mathcal{B}\circ\mathfrak{s}^{(\infty)}\circ\mathcal{B}_{\mathfrak{s}}^{-1}\right)(x^{\prime})$.
@@
$\begin{array}[]{ll}\left(\mathcal{B}\circ\mathfrak{s}^{(\infty)}\circ\mathcal{B}_{\mathfrak{s}}^{-1}\right)^{*}\left(dx^{\prime}_{1}\wedge...\wedge dx^{\prime}_{n}\right)&=\left(\mathcal{B}_{\mathfrak{s}}^{-1}\right)^{*}\circ s^{(\infty)\,*}\circ\mathcal{B}^{*}\left(dx^{\prime}_{1}\wedge...\wedge dx^{\prime}_{n}\right)\vspace{5pt}\\ &=\left(\mathcal{B}_{\mathfrak{s}}^{-1}\right)^{*}\circ s^{(\infty)\,*}\circ\left(d\left(\xi^{1}(x,u^{(k)})\right)\wedge...\wedge d\left(\xi^{n}\left(x,u^{(k)}\right)\right)\right)\vspace{5pt}\\ &=\left(\mathcal{B}_{\mathfrak{s}}^{-1}\right)^{*}\circ\left(d\left(\xi^{1}\circ s^{(k)}(x)\right)\wedge...\wedge d\left(\xi^{n}\circ s^{(k)}(x)\right)\right)\vspace{5pt}\\ &=dx_{1}\wedge...\wedge dx_{n},\end{array}$
@@
$\left\{\begin{array}[]{l}x^{\prime}_{i}=\xi^{i}\left(x,u^{(k)}\right),\vspace{10pt}\\ u^{\prime}{}^{j}=\nu^{j}\left(x,u^{(k)}\right),\end{array}\right.$
@@
$\begin{array}[]{ll}\mathcal{B}^{*}\left(\omega^{\prime}{}_{\sigma}^{j}\right)&=\mathcal{B}^{*}\left(du^{\prime}{}_{\sigma}^{j}-\sum_{i}u^{\prime}{}_{\sigma+1_{i}}^{j}dx^{\prime}_{i}\right)=d\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d\xi^{i}\vspace{5pt}\\ &=d_{H}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d_{H}\xi^{i}+d_{V}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d_{V}\xi^{i}\vspace{5pt}\\ &=\sum_{s}\left(D_{s}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}D_{s}\xi^{i}\right)dx_{s}\mod{\mathcal{I}_{\mathcal{C}(\pi)},}\end{array}$
in view of the identity $d_{V}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d_{V}\xi^{i}=0\mod{\mathcal{I}_{\mathcal{C}(\pi)}}$. Thus $\mathcal{B}^{*}\left(\mathcal{I}_{\mathcal{C}(\pi^{\prime})}\right)\subset\mathcal{I}_{\mathcal{C}(\pi)}$ if and only if $D_{s}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}D_{s}\xi^{i}=0$, that is

### v5

@@ 2.1 Differential equations as submanifolds of jet spaces
This is a very important notion in the theory of integrable equations with $2$ independent variables [1, 12, 21, 44, 47]. In such a case, by taking $\alpha=A\,dx_{1}+Bdx_{2}$, condition (1) reads $D_{2}A-D_{1}B+[A,B]=0$ on $\mathcal{E}^{(\infty)}$.
Here we notice that (1) can also be equivalently written as $\bar{d}_{H}\alpha-\alpha\wedge\alpha=0$. Moreover, since $\mathcal{E}^{(\infty)}\subset J^{\infty}(\pi)$, any element of $\mathfrak{g}\otimes\bar{\Lambda}^{1}(\mathcal{E})$ can be identified with an element of $\mathfrak{g}\otimes\bar{\Lambda}^{1}(\pi)$. Hence, in the outer geometry, (1) can also be rewritten as $d_{H}\alpha-\frac{1}{2}\left[\alpha,\alpha\right]=0\;\mbox{mod}\,\mathcal{E}^{(\infty)}$. In particular, when $\mathcal{E}=\{F^{j}=0,\,j=1,...,h\}$, under regularity assumptions (i.e., if any prolongation $\mathcal{E}^{(\infty)}$, $h\geq 0$, is totally non-degenerating [37]) equation (1) can also be rewritten in “characteristic” form $d_{H}\alpha-\frac{1}{2}\left[\alpha,\alpha\right]=\sum D_{\sigma}(F^{j})\gamma_{j}^{\sigma}$, where $\gamma_{j}^{\sigma}\in\mathfrak{g}\otimes\bar{\Lambda}^{2}(\pi)$. Hence in general (1) holds modulo differential consequences of $\mathcal{E}$, thus (1) is not equivalent to $\{F^{j}=0,\,j=1,...,h\}$.
@@ 2.2 $\mathcal{C}$-morphisms (or Lie-Bäcklund maps)
A $\mathcal{C}$-morphism, also referred to as a Lie-Bäcklund transformations [4, 25], is a smooth map $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$, such that $\mathcal{B}_{*}\mathcal{C}_{\theta}(\pi)\subseteq\mathcal{C}_{\mathcal{B}(\theta)}(\pi^{\prime})$, for any $\theta\in J^{\infty}(\pi)$, i.e.,
$\mathcal{B}_{*}\mathcal{C}(\pi)\subseteq\mathcal{C}(\pi^{\prime}).$ (2)
Notice that $\mathcal{B}_{*}\mathcal{C}(\pi)\subseteq\mathcal{C}(\pi^{\prime})$ is equivalent to $\mathcal{B}^{*}\left(\mathcal{I}_{\mathcal{C}(\pi^{\prime})}\right)\subseteq\mathcal{I}_{\mathcal{C}(\pi)}$, where $\mathcal{I}_{\mathcal{C}(\pi)}$ is the EDS generated by Cartan (or multi-contact) forms $\omega_{\sigma}^{j}=du_{\sigma}^{j}-\sum_{i}u_{\sigma+1_{i}}^{j}dx_{i}$ on $J^{\infty}(\pi)$, analogously $\mathcal{I}_{\mathcal{C}(\pi^{\prime})}$.
Notice that in general these transformations are generalizations of Lie transformations (i.e., point and contact transformations) that need not to be diffeomorphisms. Indeed, as shown by Bäcklund [6] (see also [4]), only Lie transformations are invertible. Moreover, by $\mathcal{B}^{*}\left(\mathcal{I}_{\mathcal{C}(\pi^{\prime})}\right)\subseteq\mathcal{I}_{\mathcal{C}(\pi)}$ it follows that under such a transformation the image of an integral manifolds of Cartan distribution is still an integral manifold.
@@
$\left\{\begin{array}[]{l}x^{\prime}_{i}=\xi^{i}\left(x,u^{(k)}\right),\vskip 10.0pt\\ u^{\prime}{}^{j}=\nu^{j}\left(x,u^{(k)}\right),\vskip 10.0pt\\ \vdots\\ u^{\prime}{}_{\sigma}^{j}=\nu_{\sigma}^{j}\left(x,u^{(k+|\sigma|)}\right),\qquad|\sigma|\geq 0\end{array}\right$. (3)
where in view of $\mathcal{B}^{*}\left(\mathcal{I}_{\mathcal{C}(\pi^{\prime})}\right)\subseteq\mathcal{I}_{\mathcal{C}(\pi)}$ the functions $\xi^{i}$ and $\nu_{\sigma}^{j}$ are such that $d\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d\xi^{i}=0\mod{\mathcal{I}_{\mathcal{C}}}$.
Like in the finite order case, integral manifolds $\Sigma$ of $\mathcal{C}(\pi)$ with independence condition $\text{$\Omega$=}dx_{1}\wedge...\wedge dx_{n}\neq 0$ (i.e., such that $\Omega|_{\Sigma}\neq 0$) are $\infty$-th order prolongations of sections of $\pi$. One has the following
@@
Proof.
[…]
@@
$\begin{array}[]{ll}\left(\mathcal{B}\circ\mathfrak{s}^{(\infty)}\circ\mathcal{B}_{\mathfrak{s}}^{-1}\right)^{*}\left(dx^{\prime}_{1}\wedge...\wedge dx^{\prime}_{n}\right)&=\left(\mathcal{B}_{\mathfrak{s}}^{-1}\right)^{*}\circ s^{(\infty)\,*}\circ\mathcal{B}^{*}\left(dx^{\prime}_{1}\wedge...\wedge dx^{\prime}_{n}\right)\vskip 5.0pt\\ &=\left(\mathcal{B}_{\mathfrak{s}}^{-1}\right)^{*}\circ s^{(\infty)\,*}\circ\left(d\left(\xi^{1}(x,u^{(k)})\right)\wedge...\wedge d\left(\xi^{n}\left(x,u^{(k)}\right)\right)\right)\vskip 5.0pt\\ &=\left(\mathcal{B}_{\mathfrak{s}}^{-1}\right)^{*}\circ\left(d\left(\xi^{1}\circ s^{(k)}(x)\right)\wedge...\wedge d\left(\xi^{n}\circ s^{(k)}(x)\right)\right)\vskip 5.0pt\\ &=dx_{1}\wedge...\wedge dx_{n},\end{array}$
@@
$\left\{\begin{array}[]{l}x^{\prime}_{i}=\xi^{i}\left(x,u^{(k)}\right),\vskip 10.0pt\\ u^{\prime}{}^{j}=\nu^{j}\left(x,u^{(k)}\right),\end{array}\right$.
@@
$\begin{array}[]{ll}\mathcal{B}^{*}\left(\omega^{\prime}{}_{\sigma}^{j}\right)&=\mathcal{B}^{*}\left(du^{\prime}{}_{\sigma}^{j}-\sum_{i}u^{\prime}{}_{\sigma+1_{i}}^{j}dx^{\prime}_{i}\right)=d\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d\xi^{i}\vskip 5.0pt\\ &=d_{H}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d_{H}\xi^{i}+d_{V}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d_{V}\xi^{i}\vskip 5.0pt\\ &=\sum_{s}\left(D_{s}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}D_{s}\xi^{i}\right)dx_{s}\mod{\mathcal{I}_{\mathcal{C}(\pi)},}\end{array}$
in view of the identity $d_{V}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}d_{V}\xi^{i}=0\mod{\mathcal{I}_{\mathcal{C}(\pi)}}$. Thus $\mathcal{B}^{*}\left(\mathcal{I}_{\mathcal{C}(\pi^{\prime})}\right)\subseteq\mathcal{I}_{\mathcal{C}(\pi)}$ if and only if $D_{s}\nu_{\sigma}^{j}-\sum_{i}\nu_{\sigma+1_{i}}^{j}D_{s}\xi^{i}=0$, that is